### v2

# Robust Visual Representation Learning With Multi-modal Prior Knowledge For Image Classification Under Distribution Shift

Abstract
Despite the remarkable success of deep neural networks (DNNs) in computer vision, they fail to remain high-performing when facing distribution shifts between training and testing data. In this paper, we propose Knowledge-Guided Visual representation learning (KGV) – a distribution-based learning approach leveraging multi-modal prior knowledge – to improve generalization under distribution shift. It integrates knowledge from two distinct modalities: 1) a knowledge graph (KG) with hierarchical and association relationships; and 2) generated synthetic images of visual elements semantically represented in the KG. The respective embeddings are generated from the given modalities in a common latent space, i.e., visual embeddings from original and synthetic images as well as knowledge graph embeddings (KGEs). These embeddings are aligned via a novel variant of translation-based KGE methods, where the node and relation embeddings of the KG are modeled as Gaussian distributions and translations, respectively. We claim that incorporating multi-model prior knowledge enables more regularized learning of image representations. Thus, the models are able to better generalize across different data distributions. We evaluate KGV on different image classification tasks with major or minor distribution shifts, namely road sign classification across datasets from Germany, China, and Russia, image classification with the mini-ImageNet dataset and its variants, as well as the DVM-CAR dataset. The results demonstrate that KGV consistently exhibits higher accuracy and data efficiency across all experiments.
Refer to captionRoad Sign Domain(a)Refer to captionGermany:(Train)China:(Test)Russia:(Test)Refer to captionCar DomainRefer to captionTrain:Test:Refer to captionImageNet DomainRefer to captionImageNet:(Train)ImageNet-V2:(Test)ImageNet-R:(Test)Refer to captionRefer to captionDE(b)TriangleExclamationWhiteRedhasShapehasSignLegendhasBackgroundC.hasBorderC.Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionCNDanger SignhasShapehasSignLegendBlackYellowRefer to captionRefer to captionhasBackgroundC.Refer to captionRefer to captionhasBorderC.Refer to captionRefer to captionSilverhasColorRefer to captionRefer to captionCoupeRefer to captionRefer to captionhasBodytypehasBodytypeAutohasGearType45hasViewpointNumber2hasDoorNumberhasDoorN.135hasViewpointManuellhasGearTypeThings(c)FoodBeverageBeerRefer to captionRefer to captionEatCornHotdogRefer to captionRefer to captionPlaceNatureCliff Coral Reef SpeciesEukar.AnimalBirdRefer to captionRefer to captionMammaFungus
Figure 1: (a) Examples of distribution shift in road sign, ImageNet, and Car Domains. (b) Decomposition of elements w.r.t. the road signs and cars. It represents the association relations between object categories and object category elements. (c) Examples of the abstract representation of hierarchical relations in the ImageNet domain.
Introduction
Deep neural networks (DNNs) have demonstrated outstanding performance in computer vision (CV) tasks. They achieved remarkable success by leveraging vast amounts of data and computational power. However, these models often struggle with generalization, especially when the test data distribution deviates from the training data distribution (Liang, He, and Tan 2023). This happens as machine learning theory assumes that training and test data are identically distributed. Therefore, DNNs tend to overfit to the training domain. The situation worsens under low data regimes (Adadi 2021), as DNNs are more prone to overfitting when insufficient training data exists. Figure 1a) illustrates three distinct examples of distribution shifts between train and test data, in particular the Road Sign, ImageNet and Car domain. A DNN-based classifier applied to the testing domain will deteriorate in accuracy. This happens because the visual elements of the images vary between the train and test domains. For example, in Germany, warning signs have a white background, whereas in China, they use a yellow background. Similarly, ImageNet and Car differentiate between train and test data because of their style.
Over the years, numerous approaches have been presented to utilize prior knowledge to enhance the performance of DNNs. For instance, in the field of zero-shot learning (ZSL) (Chen et al. 2023), Wang, Ye, and Gupta (2018); Kampffmeyer et al. (2019) leverage additional information from a Knowledge Graphs (KG) to classify images of unseen classes. Li et al. (2023b) combines a knowledge-based classifier and a vision-based classifier to enhance the model’s ability to classify new categories. These methods are excepted to classify unseen classes based on the other seen classes sharing the same superclass. A concrete example is the classification of unseen zebras based on seen horses and donkeys as they both belong to the same superclass, i.e., equus. The prior knowledge contained in their KG pertains to the hierarchical levels of biological classification.
However, approaches that only use symbolic knowledge from the KG are limited to addressing the issue of data distribution shift as the symbolic knowledge lacks the connection to the visual domain. We hypothesize that the model’s ability to generalize can be further improved by leveraging multi-modal prior knowledge that (partially) covers new test data distributions. The multi-model prior knowledge here indicates both a KG with hierarchical and association relations and generated synthetic images of visual elements semantically represented in the KG. To investigate this hypothesis, in this paper, we propose a novel neuro-symbolic approach, which aligns the image embeddings and knowledge graph embeddings (KGEs) in a common latent space by means of a variant of translation-based KGE method (Bordes et al. 2013). Thus, it regularizes the latent space with the guidance of the multi-modal prior knowledge to prevent the model from overfitting.
First, we capture information about the structures of objects in a KG (Hogan et al. 2021) to inform and enhance the learning process. For instance, warning signs are composed of a white triangle with a red border plus one pictographic element (cf. Figure 1b). Road signs in different countries share the same elements, such as the shape or the pictographic element, or both. Also, the hierachical relations are demonstrated in Figure 1c. Combining this information, we construct a KG and develop a translation-based KGE method to map the nodes of the KG to embeddings in the latent space. An object category (e.g., ‘danger sign’) is naturally represented by its set of image vectors in the latent space. To align image and object categories, we represent the latter by a Gaussian distribution in the common latent space. Thus, the embeddings of object category elements and their relations to object categories indirectly influence the formation of categories in the latent space.
Second, for each object category element, we generate their respective synthetic images (if applicable) to augment the dataset. For instance, in the road sign domain, these elements are shape, color, and legend within the sign. These images are considered as visual prior knowledge of the object categories defined in the KG. Therefore, in this paper, the prior knowledge we used has two distinct modalities: 1) the factual knowledge across different data distributions represented in a KG, and 2) generated synthetic images of object category elements aligned with their semantic representation in the KG.
We evaluate the performance of our approach under distribution shifts and low data regimes in the context of image classification. To represent the prior knowledge, we construct three targeted KGs for the domains of Road Sign Recognition, Car Recognition, and ImageNet. Our experimental results show that KGV outperforms the image classification baselines in all experiments. Additionally, we found that our method is more data-efficient. Our main contributions are summarized as follows:
  • We propose our KGV method, an end-to-end neuro-symbolic approach that uses multi-modal prior knowledge to regularize the latent space and thus enhance the generalization ability under data distribution shift.
  • We developed a new variant of translation-based KGE to align the embeddings of images and the KG in a common embedding space, where the nodes and predicates in the KG are represented as Gaussian distributions and translations, respectively.
  • We evaluate KGV on various image classification datasets, namely road sign datasets, mini-ImageNet and its variants, as well as the DVM-CAR datasets. The results illustrate a notable improvement under distribution shift: 4.4% and 4.1% on average in the ImageNet and Road Sign Recognition domains. Moreover, our method improves the SOTA results on the DVM-CAR dataset and shows significant improvements on data efficiency in low data regimes.
  • We analyse whether KGV can be adapted to the SOTA vision foundation models like CLIP and DINOv2. The results indicate that integrating KGV, CLIP and DINOv2 exhibit a better performance across various datasets. The code base is available at https://anonymous.4open.science/r/kgv-D68F/.
Preliminaries
Visual Embeddings
A visual encoder network $f_{\boldsymbol{\phi}}(\cdot):\mathcal{I}\rightarrow\mathbb{R}^{d}$ with parameters of $\boldsymbol{\phi}$ maps images in image space $\mathcal{I}$ to a visual embeddings $\boldsymbol{z^{I}}\in\mathbb{R}^{d}$, where $d$ is the dimension of the latent space. This function is typically realized through a series of transformations involving neural networks, particularly convolution-based models (He et al. 2015a), or Transformer-based models (Dosovitskiy et al. 2021; Carion et al. 2020).
Knowledge Graph Embeddings
[…]
Formally, we define a KG as $\mathcal{G}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ over a set $\mathcal{E}$ of entities and a set $\mathcal{R}$ of predicates. A KGE model maps entities and predicates into $d$-dimensional embedding $M_{\theta}:\mathcal{E}\cup\mathcal{R}\rightarrow\mathbb{R}^{d}$. We denote $\langle h,r,t\rangle$ as a triple in $\mathcal{G}$. A score function $s(\textbf{h},\textbf{r},\textbf{t})$ is defined on the embeddings of a triple, which should assign higher scores to positive triples, i.e., true facts in the KG, and lower scores to negative triples, i.e., corrupted triples, generated by randomly replacing the head or tail entity in an observed triple with a random entity sampled from $\mathcal{E}$. The parameter $\theta$ is learned by minimizing cross-entropy loss (Bordes et al. 2013) or margin-based loss (Sun et al. 2019).
Few-shot Learning
A classifier $f:x\rightarrow y$ is initially trained with a set of labeled training samples $\mathcal{D}_{\textrm{tr}}=\{(x,y)|x\in\mathcal{X},y\in\mathcal{Y}\}$. Few-shot learning focuses on a common scenario called $N$-way, $K$-shot learning. Given a set of $K$ labeled examples for each of $N$ different classes $\mathcal{D}_{\textrm{few}}=\{(x,y)|x\in\mathcal{X}_{\textrm{few}}^{\prime},y\in\mathcal{Y^{\prime}}\}$, where $|\mathcal{Y^{\prime}}|=N$ and $\forall y_{0}\in\mathcal{Y^{\prime}},|\{(x,y)\in\mathcal{D}_{\textrm{few}}|y=y_{0}\}|=K$, in the target dataset, this classifier is able to correctly predict the labels of samples in the full target dataset $\mathcal{D}_{\textrm{full}}=\{(x,y)|x\in\mathcal{X}_{\textrm{full}}^{\prime},y\in\mathcal{Y^{\prime}}\}$, with $(\mathcal{X}\cup\mathcal{X}_{\textrm{few}}^{\prime})\cap\mathcal{X}_{\textrm{full}}^{\prime}=\emptyset$.
Related Work
Use Prior Knowledge for Learning Tasks
Implicit prior knowledge in pre-trained models like CLIP (Radford et al. 2021) and DINOv2 (Oquab et al. 2024) can be leveraged by continuing to fine-tune them. Including explicit knowledge from KGs can also enhance model’s ability in visual or textual tasks (Monka, Halilaj, and Rettinger 2022b). Some approaches (Annervaz, Chowdhury, and Dukkipati 2018; Bosselut et al. 2019; Liu et al. 2020) leverage KGs to enhance language models with better reasoning ability. In the field of computer vision, KGs also play an important role. Li et al. (2023a) introduce a KG that includes inter-class relationships to the visual-language model, yielding a more effective classifier for downstream tasks. Some approaches  (Zareian, Karaman, and Chang 2020; Yu et al. 2021) build scene graphs to enhance their models’ performance. There also exist some approaches use extra information in the KG to perform zero-shot learning tasks (Wang, Ye, and Gupta 2018; Kampffmeyer et al. 2019; Roy et al. 2020) or few-shot learning tasks (Li et al. 2023c; Tseng et al. 2020).
In the field of symbolic AI, some approaches include textual information (Nayyeri et al. 2023), or multi-modal information (Chen et al. 2022; Liang et al. 2023), to increase the knowledge graph reasoning ability. They combine the structured information from KGs, visual information, and textual information with transformer-based approaches. These works prove that including visual or textual information or both is helpful for knowledge graph based reasoning and knowledge graph completion.
Integrate Knowledge Graphs with Other Modalities
Two primary approaches are recognized in the literature to encapsulate the methodologies for integrating heterogeneous information from KGs with other modalities, such as visual and textual information. The first category (Chen et al. 2022; Liang et al. 2023; Mousselly-Sergieh et al. 2018) involves merging structural data derived from KGs — typically obtained via Graph Neural Networks (GNNs)—through the utilization of cross-attention mechanisms as employed in Transformer models. The second category (Li et al. 2023a; Monka et al. 2021; Monka, Halilaj, and Rettinger 2022a) entails aligning KGEs with visual or textual embeddings in one common space by contrastive learning (Radford et al. 2021).
In this paper, we propose a novel approach that leverages the multi-modal prior knowledge from both, a KG and synthetic images generated based on visual features that are semantically represented in the KG, in order to boost the model’s ability to generalize under data distribution shifts. Our approach aligns KGEs and visual embeddings in the same latent space using a new variant of translation-based KGE methods, where the nodes and relations are represented as Gaussian distributions and translations, respectively, to model the structured information in the KG within the latent space. Existing translation-based KGE methods typically represent entities and predicates with vector embeddings (Bordes et al. 2013; Wang et al. 2014; Lin et al. 2015; Socher et al. 2013). The authors (He et al. 2015b; Xiao et al. 2015) take the uncertainty of entities and predicates into account and represent them by probability distributions.
Methodology
Overview
We propose KGV, an approach to address the limitations of the deep learning solutions for supervised image classification and few-shot learning w.r.t. to the distribution shifts. The core idea of KGV is the inclusion of multi-modal prior knowledge coming from two different modalities: 1) domain knowledge - captured in a semantic KG; and 2) synthetic images - generated for object category elements. We encode the node and relation information from the KG and images as KGEs ($\boldsymbol{r},\boldsymbol{t}$) and image embeddings $\boldsymbol{z^{I}}$ into one latent space and align them via a variant of translation-based KGE method. Here, we represent the object categories in KG as Gaussian embeddings in the latent space, and we assume each image embedding should follow the distribution of its corresponding category distribution. It extends the contrastive learning, as seen in SupCon (Khosla et al. 2020) and CLIP (Radford et al. 2021), which focuses solely on similarity and dissimilarity, by incorporating the capability to model additional individual relationships between entities.
Our approach is designed end-to-end by adding the regularization loss and cross-entropy loss together to train the network. The regularization loss aligns the KG embedding and the visual embedding. The cross-entropy loss is used for the image classification. The architecture of KGV is depicted in Figure 2. Our model consists of three phases: the knowledge modeling phase, the training phase, and the inference phase. Here, we take the road sign domain as an example to describe our model.
Knowledge Modeling
Prior knowledge used in our KGV exists in two distinct modalities. Firstly, an expert-constructed KG includes factual knowledge such as hierarchical and association relations across different data distributions. Secondly, we generate synthetic images of object category elements that are semantically represented in the KG but lack visual information in the image dataset. In the road sign recognition domain, these elements are shape, color, and the sign legend.
Knowledge Graph Construction:
1. Knowledge Modelling PhaseRoad Sign DomainColorsRefer to captionRefer to captionRefer to captionShapesRefer to captionRefer to captionRefer to captionSign LegendsRefer to captionRefer to captionRefer to captionRoad SignsRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionVisual Knowledge GraphRoad SignWarningMandatory Danger
Animal
Danger
Sign
[…]
$\hat{y}_{2}$ $\hat{y}_{1}$ $\hat{y}_{N_{\textrm{cls}}}$
Figure 2: The KGV architecture – Our approach consists of three phases, namely knowledge modeling, training, and inference. Knowledge modeling phase: We create a knowledge graph based on prior domain knowledge. Also, synthetic images are generated for object categories (e.g., shapes and colors) that are semantically represented in the knowledge graph but lack visual information in the dataset. Training phase: The neural network is fed with both synthetic and real-world images and trained end-to-end by adding the regularization loss and cross-entropy loss together as a total loss for optimization. The image embeddings $\boldsymbol{z^{I}}$ and knowledge graph embeddings $\boldsymbol{z^{r}_{i}}$, $\boldsymbol{z^{o}_{j}}$ are aligned by minimizing the regularization loss. The Cross-entropy loss is used to classify images based on their image embedding representation. Inference phase: The classification task is completed by selecting the class with the highest possibility based on the output of the decoder.
We construct a KG to represent object categories, object category elements, and the relationships between an object category and its elements based on domain knowledge. At the top level, there are two different categories: road sign and road sign feature. The road sign category comprises more sub-categories: informative, prohibitory, mandatory, and warning. Further, we define dedicated classes, such as Germany, China, and Russia, to model various country versions of the road signs. The road sign feature category contains sub-categories of shape, color, and legend within the sign. An illustration of hierarchical relations is demonstrated in the appendix. Besides the hierarchical information, the KG also contains four association relations. They are ‘has the shape of’, ‘has the sign legend of’, ‘has the background color of’, and ‘has the border color of’. For instance, as Figure 1b shows, the danger sign has the shape of a triangle and has the sign legend of the exclamation mark.
Synthetic images of object category elements:
We generate synthetic images of object category elements such as shapes and colors, which is treated as visual prior knowledge. In our setting, images of different colors are generated by randomly selecting an image within the color range corresponding to the popular color definition. The images of shapes are generated by alternating on the thickness, size, and location on the canvas. The images of sign legends are extracted from standard road signs, then randomly resized and positioned at various locations on the canvas. These synthetic images are then added to the road sign dataset and trained together with those road sign images. All types of synthetic images can be seen in the appendix.
Training
Our training loss comprises two components: the regularization loss, which is used to align the image embeddings with the KGEs, and the cross-entropy loss, which is employed for classification purposes. The regularization loss is specifically formulated to maximize the score function value for positive triplets and minimize it for negative triplets. In the following subsections, we first delineate the image, node, and relation embeddings. Subsequently, we elucidate the methodology for constructing triplets of embeddings and the process for masking the positive and negative triplets. Next, we introduce the design of the score function for triplets. Finally, we provide a summary of the overall loss function.
Embeddings:
Each image is provided as input into the image encoder to obtain the respective image embedding $\boldsymbol{z^{I}}\in\mathbb{R}^{d}$ in the latent space, where $d$ is the size of the dimension. The relations from KG connecting object categories, their elements and their respective images are encoded and stored in the relation lookup table, which is represented as a list of embeddings $[\boldsymbol{z_{0}^{r}},\boldsymbol{z_{1}^{r}},...,\boldsymbol{z_{N_{r}}^{r}}]$, where each $\boldsymbol{z_{i}^{r}}$ is the $i$-th relation vector embedding with the size of $d$. Note that the 0-th relation represents the inclusion relation demonstrated with purple dashed arrows as illustrated in Figure 2. The hierarchy of road signs is preserved in the KG. For instance, the images of the danger signs in Germany belong to the German danger sign category, and they also belong to the danger sign category, warning sign category, and road sign category. We use $\boldsymbol{z_{0}^{r}}$ to represent all these inclusion relations. In terms of the other $N_{r}$ relations, if one node $h$ has one relation with another node $t$, then all the sub-category nodes of $h$ have such relation with the node $t$. For example, all warning signs have a triangle shape, including their sub-categories like danger and danger animal. The node lookup table is represented as a list of embeddings $[\boldsymbol{z_{1}^{o}},\boldsymbol{z_{2}^{o}},...,\boldsymbol{z_{N_{o}}^{o}}]$, where each $\boldsymbol{z_{j}^{o}}$ denotes the $j$-th embedding. $N_{o}$ is the total number of nodes in the KG. The node embeddings are considered as Gaussian embeddings $\boldsymbol{z_{j}^{o}}=(\boldsymbol{\mu_{j}},\boldsymbol{\Sigma_{j}})$, where $\boldsymbol{\mu_{j}}$ and $\boldsymbol{\Sigma_{j}}$ are the mean and variance of the Gaussian distribution. Note that nodes, their relations, and respective images are embedded in the same latent space.
Positive and Negative Triplets of Embeddings:
Following the close world assumption in the context of KG, any triplet not explicitly defined in the KG is considered to be a negative one. Since $N_{o}$ nodes and $N_{r}+1$ relations exist in the KG, for each image embedding $\boldsymbol{z^{I}}$, $(N_{r}+1)\times N_{o}$ triplets with the embedding form of $\langle\boldsymbol{z^{I}},\boldsymbol{z^{r}_{i}},\boldsymbol{z^{o}_{j}}\rangle$ can be created. The positive triplets represent the defined information in the KG about the given image. To mask all the positive triplets and negative triplets for one given image, we construct a tensor $\boldsymbol{M}\in[0,1]^{(N_{r}+1)\times N_{o}}$:
$\boldsymbol{M}_{ij}=\begin{cases}1&\text{if }\langle e_{node(y)},r_{i},e_{j}\rangle\in\mathcal{G}\\ 1&\text{if }r_{i}=\text{instanceOf}\wedge j=node(y)\\ 0&\text{otherwise}\end{cases}$ (1)
where $e_{j}$ represents the $j$-th element in entity set $\mathcal{E}$. $r_{i}$ is the $i$-th element in the relation set $\mathcal{R}$. $\mathcal{G}$ represents the triplet stored in the KG. $y$ is the class label of the given image. $node(\cdot)$ indicates the mapping function connecting the class id of the image to the corresponding node id in the KG. As Figure 2 shows, positive triplets are marked with green color, whereas negative triplets are marked with white color.
Score Function Design:
For a given triplet $\langle\boldsymbol{z^{I}},\boldsymbol{z_{i}^{r}},\boldsymbol{z_{j}^{o}}\rangle$, we design the score function in translation-based manner which means $\boldsymbol{z^{I}}+\boldsymbol{z_{i}^{r}}\approx\boldsymbol{z_{j}^{o}}$. In this paper, we represent image embeddings and relation embeddings as vector embeddings. The node embeddings are considered as Gaussian embeddings $\boldsymbol{z_{j}^{o}}=(\boldsymbol{\mu_{j}},\boldsymbol{\Sigma_{j}})$ to better handle the inclusion relationship between the image embedding and the node embedding, where $\boldsymbol{\mu_{j}}$ and $\boldsymbol{\Sigma_{j}}$ are the mean and variance of the Gaussian distribution. The intuition behind this is that each node in the KG contains various image instances. Thus, it is more reasonable to model the node embedding as Gaussian since each node represents a cluster of image vector embeddings. The score function can be defined as:
$s(\boldsymbol{z^{I}},\boldsymbol{z_{i}^{r}},\boldsymbol{z_{j}^{o}})=\mathcal{N}(\boldsymbol{z^{I}}+\boldsymbol{z_{i}^{r}};\boldsymbol{\mu_{j}},\boldsymbol{\Sigma_{j}})$ (2)
where $\mathcal{N}(\boldsymbol{z^{I}}+\boldsymbol{z_{i}^{r}};\boldsymbol{\mu_{j}},\boldsymbol{\Sigma_{j}})$ denotes the possibility density of the vector $\boldsymbol{z^{I}}+\boldsymbol{z_{i}^{r}}$ under the Gaussian distribution with parameters of $\boldsymbol{\mu_{j}}$ and $\boldsymbol{\Sigma_{j}}$. Different score function designs are also compared in our experiments.
Loss Functions:
During the training phase, there are two losses to minimize. The first one is the cross-entropy loss $\mathcal{L}_{\textrm{CE}}(\boldsymbol{\hat{y}},\boldsymbol{y})$ to classify the images to their corresponding classes, where $\boldsymbol{\hat{y}}$ is the predicted possibility of classes and $\boldsymbol{y}$ is the ground truth class label in the form of one-hot-key.
The other loss term is the regularization loss. It aims to regularize the latent space with prior knowledge by aligning image embeddings with KG embeddings. Here, we represent the relationships from the KG in the visual latent space by optimizing the score function as defined in Equation (2). If the score function has high scores for positive triplets and low scores for negative triplets, the image embeddings and the KGEs are then well aligned in the latent space. The regularization loss is defined as follows:
$\mathcal{L}_{\textrm{reg}}=\frac{\sum_{i,j}\boldsymbol{M}_{ij}\cdot\boldsymbol{S}_{ij}}{\sum_{i,j}\boldsymbol{M}_{ij}}+\frac{\sum_{i,j}(1-\boldsymbol{M}_{ij})\cdot max\{0,\epsilon-\boldsymbol{S}_{ij}\}}{\sum_{i,j}1-\boldsymbol{M}_{ij}}$ (3)
where M is the mask defined in Equation (1). $\textbf{S}\in R^{(N_{r}+1)\times N_{o}}$ and $\textbf{S}_{ij}=-\log(s(\boldsymbol{z^{I}},\boldsymbol{z_{i}^{r}},\boldsymbol{z_{j}^{o}}))$, following the negative log likelihood form. $\epsilon$ is the threshold score difference between positive and negative triplets. Finally, the total loss is defined by:
$\mathcal{L}=\mathcal{L}_{\textrm{CE}}+\beta\mathcal{L}_{\textrm{reg}}\text{,}$ (4)
where $\beta$ is the hyper-parameter used to balance the cross-entropy loss term and regularization loss term.
Inference Phase
At the inference phase, the classification task is completed by selecting the class with the highest probability based on the output of the decoder $\boldsymbol{\hat{y}}$.
Experiments
In this section, we first explain the datasets and the respective knowledge graphs used in our experiments. Then, we analyze the results of supervised image classification experiments under varying conditions, including data distribution shifts, low-data regimes, and few-shot learning scenarios. The experimental details are listed in the appendix section.
Datasets
In our experiments, we evaluate KGV in three domains, namely Road Sign, Car Recognition, and ImageNet.
  • Road Sign Recognition: It includes three commonly used road sign datasets from Germany, China, and Russia: German Traffic Sign Recognition Benchmark (GTSRB) from Stallkamp et al. (2011), Chinese Traffic Sign Dataset (CTSD) from Yang et al. (2015), and Russian Traffic Sign Images Dataset (RTSD) from Shakhuro and Konouchine (2016).
  • Car Recognition: We use the Deep Visual Marketing Car (DVM-CAR) dataset  (Huang et al. 2021) which contains millions of car images captured under various conditions and viewpoints, accompanied by detailed annotations including car make, seat count, door count, color, body type and etc. The primary task here is to classify the car model based on the provided image.
  • ImageNet: In this domain, we use the well-known mini-ImageNet (Vinyals et al. 2016), ImageNet-Sketch (Wang et al. 2019), ImageNetV2 (Recht et al. 2019), ImageNet-R (Hendrycks et al. 2021a), and ImageNet-A (Hendrycks et al. 2021b) for our experiments.
Knowledge Graphs
Road Sign Recognition Knowledge Graph:
The constructed knowledge graph contains 202 nodes and 5 relations. The 5 relations we include are ‘has the background color of’, ‘has the border color of’, ‘has the sign legend of’, ‘has the shape of’, ‘instance/subclass’.
Car Recognition Knowledge Graph:
The knowledge graph for the DVM-CAR contains 386 nodes and 9 relations. These relations are ‘hasBodyType’, ‘hasColor’, ‘hasDoorNo’, ‘hasFuelType’, ‘hasGearbox’, ‘hasMaker’, ‘hasSeatNo’, ‘hasViewpoint’, ‘instanceOf’. The dataset includes annotations such as body type, color, and car make for each image, provided in tabular form. We construct a KG by defining the ontology and triplets based on the tabular annotations. We generate pure color images using the same method as in the road sign domain to serve as visual prior knowledge. Further, we collect sketch images representing body types such as SUV, Coupe, and Limousine. Examples of these body type sketches are provided in the appendix.
ImageNet Knowledge Graph:
We leverage WordNet (Miller 1995; Fellbaum 2010), which is a lexical database containing nouns, verbs, adjectives, and adverbs of the English language structured into respective synsets and create the subset of Wordnet by extracting the respective synsets of each label from the mini-ImageNet dataset. These synsets are grouped based on the lexical domain they pertain to, e.g., animal, artifact, or food; thus, we get hierarchical relations between the classes in the mini-ImageNet from the WordNet. We also include two association relations: ‘has color’ and ‘has sketch’. We use the generated pure color images in the domain of road signs and the sketch images from ImageNet-Sketch. It contains 234 nodes and 3 relations: ‘subclass/instance of’, ‘has color’ and ‘has sketch’.
Results
We conducted a number of supervised image classification experiments with major and minor distribution shifts and low data regimes. Further, we evaluated the model’s adaptability on few-shot learning experiments using data with altered distributions.
Image Classification under Data Distribution Shifts:
Prior Know. Model GTSRB (43) CTSD (25) RTSD (36)
Imp. KG VP
ResNet50 98.9% (+0.4%) 71.5% (+3.3%) 72.8% (+7.3%)
CLIP+LP 77.2% (+22.1%) 73.5% (+1.3%) 60.0%(+20.1%)
DINOv2+LP 78.7% (+20.6%) 28.9% (+49.5%) 59.9% (+20.2%)
DGP 99.0 %(+0.3%) 72.0% (+2.8%) 73.5% (+6.6%)
GCNZ 99.1% (+0.2%) 71.1% (+3.7%) 74.9% (+5.2%)
KGV (ours) 99.3% 74.8% 80.1%
ResNet50+ 98.9% 71.8% 72.7%
KGV- 99.0% 72.5% 74.0%
Table 1: Road Sign Classification Accuracy. Evaluation results with the test data of GTSRB and the full data of CTSD and RTSD. Imp. refers to the implicit prior knowledge in pre-trained models, while KG indicates explicit prior knowledge from the knowledge graph. VP denotes the visual prior associated with object category elements.
Classes Images ResNet50 KGV
Precision Recall Precision Recall
STOP [Uncaptioned image] [Uncaptioned image] 21.3% 74.6% 48.2% 98.1%
Danger [Uncaptioned image] [Uncaptioned image] 0.0% 0.0% 10.8% 14.9%
BendLeft [Uncaptioned image] [Uncaptioned image] 0.0% 0.0% 3.5% 32.1%
PedestrianCross [Uncaptioned image] [Uncaptioned image] 0.0% 0.0% 9.1 % 65.6%
SchoolCross [Uncaptioned image] [Uncaptioned image] 0.0% 0.0% 20.1% 25.4%
Table 2: Precision and recall of representative classes for road sign classification when training with 43 classes data of GTSRB and testing with shared 25 classes data of CTSD.
We train our KGV model and ResNet50 with the GTSRB dataset plus generated synthetic images from scratch. It is then evaluated with test data of GTSRB with 43 classes, the full data of CTSD with 25 shared classes, and the full data of RTSD with 36 shared classes. The shared classes here indicate those classes that both datasets contain. As shown in Table 1, all models’ accuracy drops in distribution shift scenarios, i.e., from Germany to China or Russia. By incorporating the multi-modal prior knowledge, our KGV reaches an obvious improvement compared to other baselines, such as DGP (Kampffmeyer et al. 2019) and GCNZ (Wang, Ye, and Gupta 2018). These two are ZSL baselines, which leverage additional information from KGs. Fine-tuning pretrained CLIP and DINOv2 with linear probing does not work well in road sign domain. ‘ResNet50+’ refers to the model trained with additional synthetic images and ‘KGV-’ denotes the model trained without synthetic images. As the experiments indicate, our KGV model surpasses the ResNet50 baseline by 3.3% in CTSD and 7.3% in RTSD. We also observe that the inclusion of synthetic images featuring elements of object categories, such as colors and shapes, enhances the generalization ability of the KGV models across various data distributions. On the other hand, it does not have a significant effect on the ResNet50’s performance. The precision and recall for a subset of classes in CTSD are presented in Table 2. As the table illustrates, the ResNet50 fails to classify specific classes, including danger signs, pedestrian crossing, etc. In contrast, the KGV models exhibit significant improvements in both precision and recall.
Datasets Samples ResNet50 KGNN ResNet50+ KGV CLIP+LP DINOv2+LP
Mini-ImageNet [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 62.6% +4.2% 66.2% +0.6% 59.6% +7.2% 66.8% 87.0% 94.9%
ImageNetV2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 50.5% +2.2% 52.5% +0.2% 48.0% +4.7% 52.7% 87.5% 91.5%
ImageNet-R [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 25.5% +15.9% 30.0% +11.4% 36.9% +4.5% 41.4% 90.6% 86.1%
ImageNet-A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 4.7% +2.1% 5.3% +1.5% 5.0% +1.8% 6.8% 53.8% 84.5%
Table 3: Classification Accuracy in the ImageNet Domain. ResNet50+ refers to the ResNet50 model trained with additional visual prior images.
We also evaluate our method in the ImageNet domain. The baselines we choose are ResNet50, DINOv2, CLIP, and KGNN (Monka, Halilaj, and Rettinger 2022a). KGNN represents another approach to integrate prior knowledge into learning pipeline through pre-training. Our model outperforms the baselines KGNN and ResNet50 in all these datasets. As Table 3 shows, these baselines fail, particularly when the distribution shift increases (from Mini-ImageNet to ImageNetV2, ImageNet-R, and ImageNet-A). Our KGV suffers less when data distribution changes. We also observe that incorporating visual priors without the guidance of symbolic knowledge from the KG may negatively impact performance in scenarios where the difference between training and test distributions is minimal. Compared to ResNet50, ResNet50+’s performance decreases from 62.6% to 59.6% on mini-ImageNet and from 50.5% to 48.0% on ImageNetV2. We believe this is due to the visual priors altering the distribution of the training data, making it different from the test distribution. However, DINOv2 and CLIP fine-tuned with linear probing achieve better performance than our model, as they are pre-trained on massive data, including ImageNet and its variants. We conducted additional experiments by combining DINOv2 and CLIP with our approach, respectively, as described in Section Combining Implicit and Explicit Prior Knowledge.
Model SimCLR MMCL SupCon ResNet50 CLIP+LP DINOv2+LP KGV (Ours)
ACC 88.8% +7.5% 94.4% +1.9% 93.8% +2.5% 94.3% +2.0% 65.8% +30.5% 89.3% +7.0% 96.3%
Table 4: Accuracy in DVM-CAR Classification.
Next, we evaluate KGV in the DVM-CAR dataset. We observe that even though the test and training data have no significant distribution shift, it still improves performance. The baselines we choose are contrastive learning approaches such as SimCLR (Chen et al. 2020) and SupCon (Khosla et al. 2020) as well as the SOTA model MMCL (Hager, Menten, and Rueckert 2023), which combines the tabular data and image data to enhance the model’s performance through contrastive learning. Table 4 shows that our model outperforms all baselines. In the mini-ImageNet domain, KGV achieves a 4.2% performance gain compared to ResNet50.
Performance in Low Data Regimes:
Figure 3 shows the evaluation results KGV in low data regimes. We see that our method significantly outperforms the ResNet50 baseline, particularly when the amount of training data is limited. With the increase of training data, the performance of the two models gradually converges. However, KGV consistently outperforms the baseline among all training settings. The performance of KGV- – the variant of KGV trained without synthetic images – is also shown in Figure 3. It indicates that without the visual priors, its performance drops.
(a) Road Sign Classification in GTSRB.
(b) DVM-CAR Model Classification.
Figure 3: Performance under Low Data Regime. KGV- is the KGV variant trained without synthetic images.
Few-shot Learning:
Experiments Models CTSD (58 classes) RTSD (64 classes)
One-shot ResNet50 68.4% (+7.0%) 57.9% (+5.2%)
KGV 75.4% 63.1%
Five-shot ResNet50 98.7% (+0.2%) 81.2% (+2.3%)
KGV 98.9% 83.5%
Table 5: One-shot and five-shot learning results in the datasets of CTSD and RTSD are demonstrated in the table.
Prior Knowledge Model Road Sign Domain ImageNet Domain Car Domain
Imp. KG GTSRB CTSD RTSD Mini-ImageNet ImageNetV2 ImageNet-R ImageNet-A DVM-CAR
P-ResNet50+KGV 99.5% 74.5% 79.8% 81.8% 62.4% 56.6% 6.4% 97.0 %
P-ResNet50 99.3% (+0.2%) 73.1% (+1.4%) 72.5% (+7.3%) 75.0% (+6.8%) 61.3% (+1.1%) 51.6% (+5.0%) 4.4% (+2.0%) 96.2% (+0.8%)
DINOv2+KGV 79.1% 31.2% 62.4% 95.4% 92.4% 87.0% 85.4% 92.8%
DINOV2+LP 78.7% (+0.4%) 28.9% (+2.3%) 59.9% (+2.5%) 94.9% (+0.6%) 91.5% (+0.9%) 86.1% (+0.9%) 84.3% (+1.1%) 89.3% (+3.5%)
CLIP-I+KGV 89.3% 73.0% 67.8% 90.1% 89.1% 83.2% 50.1% 90.5%
CLIP-I+LP 89.1% (+0.2%) 73.5% (-0.5%) 60.0% (+7.8%) 89.7% (+0.4%) 88.9% (+0.2%) 82.9% (+0.3%) 49.6% (+0.5%) 89.2% (+1.3%)
Table 6: Pre-trained ResNet50, DINOv2 (1B), and CLIP Image encoder, which are denoted as P-ResNet50, DINOv2, and CLIP-I, are used for our experiments. LP represents linear probing for fine-tuning. Our KGV’s encoder is replaced by one of these pre-trained image encoders denoted by X+KGV.
Model Road Sign Domain ImageNet Domain Car Domain
GTSRB CTSD RTSD Mini-ImageNet ImageNetV2 ImageNet-R ImageNet-A DVM-CAR
KGV 99.3% 74.8% 80.1% 66.8% 52.7% 41.4% 6.8% 96.3%
KGV w/o Gaussian Emb. 99.4% (-0.1%) 74.7% (+0.1%) 79.9% (+0.2%) 61.2% (+5.6%) 49.0% (+3.7%) 37.8% (+3.6%) 4.8% (+2.0%) 96.3%(+0.0%)
KGV w/ TransH 99.3% (0.0%) 73.5% (+1.3%) 79.3% (+0.0%) 61.8% (+5.0%) 48.2% (+4.5%) 38.6% (+2.8%) 5.5% (+1.3%) 95.8% (+0.5%)
KGV w/ DistMult 99.2% (+0.1%) 73.7% (+1.1%) 75.6% (+5.4%) 60.5% (+6.3%) 47.6% (+5.1%) 37.5% (+3.1%) 4.7% (+2.1%) 95.1% (+1.2%)
KGV w/ RGCN 99.3% (+0.0%) 74.1% (+0.7%) 78.5% (+1.6%) 63.5% (+3.3%) 50.6% (+2.1%) 39.7% (+1.7%) 6.1% (+0.7%) -
Table 7: KGV with alternative score function designs like TransH, DistMult, and replacing Gaussian embeddings with vector embeddings are listed in the table. KGV with GNN-based method like RGCN are also compared.
We conduct our few-shot learning tasks in the domain of road sign recognition. We first train the models with the GTSRB dataset. Then, we keep the encoder network and replace the decoder to fit the right class dimension in the target dataset. The new decoder is randomly initialized. The models are retrained with different amounts of labeled target data in the CTSD or the RTSD. We perform one-shot and five-shot learning experiments. One-shot learning involves using just one image per class from the target dataset for retraining. Similarly, five-shot learning uses five images per class in the target dataset for retraining. As Table 5 shows that KGV achieves 7.0% performance gain in the CTSD dataset and 5.2% in the RTSD dataset in one-shot learning. In a five-shot learning setting, the KGV model achieves 0.2 % improvement in CTSD and 2.3% improvement in RTSD.
Combining Implicit and Explicit Prior Knowledge
In this sub-section, we investigate if combining implicit prior knowledge from the pre-trained models and explicit multi-modal prior knowledge from KG and visual prior images can further improve our method’s performance. We conduct experiments by replacing the random initialized ResNet50 encoder with pre-trained ResNet50, DINOv2(1B), and the image encoder of CLIP. As shown in Table 6, applying our method to these pre-trained encoders consistently enhances performance across all three domains. Although DINOv2 and CLIP have already achieved strong results on the ImageNet domain, the incorporation of KGV further improves their performance. These results indicate that our method can be adapted to state-of-the-art pre-trained image encoders to further enhance their performance.
Ablation Study
Score Functions
Here, we compare four score function designs: (1) replacing the Gaussian-form node embedding with vector embedding (TransE), (2) using TransH’s score function, (3) adopting DistMult’s score function, and (4) incorporating R-GCN, a GNN-based method. As Table 7 shows, overall KGV performs best among all these score function designs. The performance of KGV without Gaussian embedding drops significantly in the ImageNet domain compared to other domains. This highlights the importance of using Gaussian embedding for node representation to handle inclusion relations between object categories and their respective images. The DistMult demonstrates suboptimal performance across all three visual domains. While GNN-based method R-GCN outperform TransH and DistMult, it still does not surpass the effectiveness of our original score function design.
Knowledge types
Here, we further analyze the impact of relations types in the KGV’s performance. We conduct ablation experiments on road sign recognition, car recognition.
As Table 8 demonstrates, in each study, one type of relation is excluded, i.e. sign legends, colors, and shapes. Our findings indicate that shape is the most critical factor in enhancing performance under the distribution shift from German to Chinese and Russian road signs. This is likely because road signs with the same semantic meaning share identical shapes in Germany, China, and Russia. Sign legends also significantly contributes to performance, while information about colors is the least impactful.
Model CTSD RTSD
KGV 74.8% 80.1%
KGV w/o sign legend 73.4% (+1.4%) 75.5% (+4.6%)
KGV w/o colors 73.7% (+1.1%) 77.2% (+2.9%)
KGV w/o shapes 72.8% (+2.0%) 74.7% (+5.4%)
Table 8: Classification Accuracy in the Road Sign Domain. The models are trained with 43 classes data of GTSRB and testing with shared 25 classes data of CTSD and shared 36 classes data of RTSD.
On the DVM-CAR dataset, we performed experiments in a setting with 5% training data. We observe that even with this amount of data, KGV exhibits a large performance margin compared to the baselines. As indicated in Table 9, the information about the body type is the most important factor in improving performance. Without this information, there is a significant drop of 4.0%. Additionally, prior knowledge of colors is also crucial, as the performance drops by 2.4% without color information.
We also found that non-morphometric features, such as gear type and fuel type, can also influence performance. This is likely because specific car models are often available only with certain types of gear and fuel. For instance, Tesla primarily produces electric cars, and most of them are automatic. The prior knowledge of the image viewpoint does not play a significant role.
Model Accuracy Model Accuracy
KGV 77.1% ResNet50 64.0% (+13.1%)
KGV w/o body type 73.1% (+4.0%) KGV w/o color 74.7% (+2.4%)
KGV w/o viewpoint 76.2% (+0.9%) KGV w/o non-morpho. 75.2% (+1.9%)
Table 9: Accuracy in DVM-CAR Model Classification with 5% Training Data.
Conclusion and Future Work
In this paper, we introduce KGV, a knowledge-guided visual representation learning method that leverages multi-modal prior knowledge to improve generalization in deep learning. The multi-modal prior knowledge across data distributions in KGV helps to mitigate overfitting the training dataset, thereby enhancing the robustness of the model against data distribution shifts. We observe that modeling node embeddings of the KG as Gaussian embeddings enhances the performance of KGV to model hierachical relations. Furthermore, our method demonstrates a strong tendency to be data efficient, i.e consistently achieves good performance in the low data regimes. Finally, we show that our KGV method can be well adapted to the SOTA visual foundation models such as DINOv2 and CLIP.
A main objective of our future work is to scale our concept to larger datasets and KGs. Additionally, we plan to explore further methods for aligning KGEs and image embeddings, and inducing information of the KG at inference time. Lastly, we will look whether incorporating additional text information will improve the performance.